# On total transitivity of graphs

### v1

@@ 5 Polynomial-time algorithm for trees
In the previous section, we have seen that Maximum Total Transitivity Decision Problem is NP-complete for bipartite graphs. In this section, we design a polynomial-time algorithm for finding the total transitivity of a given tree $T=(V,E)$. First, we give a comprehensive description of our proposed algorithm.
@@ 5.1 Description of the algorithm
To this end, we define another two parameters, namely the modified-total transitive number and rooted modified-total transitive number. The modified-total transitive number of a vertex $v$ in $T$ is the maximum integer $p$ such that $v\in V_{p}$ in a modified-total transitive partition $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$. Note that the maximum is taken over all modified-total transitive partitions of $T$. We denote the modified-total transitive number of a vertex $v$ in $T$ by $mttr(v,T)$. Finally, we define another parameter, namely the rooted modified-total transitive number. The rooted modified-total transitive number of $v$ in $T^{c}$ is the modified-total transitive number of $v$ in the tree $T_{v}^{c}$, and it is denoted by $mttr^{r}(v,T^{c})$. Therefore, $mttr^{r}(v,T^{c})=mttr(v,T_{v}^{c})$.
Note that the value of the rooted total transitive number of a vertex is dependent on the rooted tree, whereas the total transitive number is independent of the rooted tree. Also, for the root vertex $c$, $ttr^{r}(c,T^{c})=ttr(c,T)$. We recursively compute the rooted total transitive number or rooted modified-total transitive number of the vertices of $T^{c}$ in a bottom-up approach. First, we consider a vertex ordering $\sigma$, the reverse of the BFS ordering of $T^{c}$. For a leaf vertex $c_{i}$, we set $mttr^{r}(c_{i},T^{c})=1$ and for a support vertex $c_{j}$, $ttr^{r}(c_{j},T^{c})=mttr^{r}(c_{j},T^{c})=1$. For other non-leaf vertex $c_{i}$, which is not a support vertex, we call the function Total_Transitive_Number()()( ), which takes the rooted total transitive number or rooted modified-total transitive number of children of $c_{i}$ in $T^{c}$ as input and returns the rooted total transitive number of $c_{i}$ in $T^{c}$ or rooted modified-total transitive number $c_{i}$ in $T^{c}$ according to Theorem 27. At the end of the bottom-up approach, we have the rooted total transitive number of $c$ in $T^{c}$ or rooted modified-total transitive number of $c$ in $T^{c}$, that is, $ttr^{r}(c,T^{c})$ or $mttr^{r}(c,T^{c})$, which is the same as the total transitive number of $c$ in $T$, that is, $ttr(c,T)$ or $ttr(c,T)+1$. After the bottom-up approach, we have the total transitive number of the root vertex $c$ and the rooted total transitive number or rooted modified-total transitive number of every other vertex in $T^{c}$. Similarly, we compute the total transitive number of all vertices of $T$ by considering a rooted tree rooted at vertex $x$. Finally, we compute the maximum overall total transitive number of vertices in $T$ and return that value as the total transitivity of $T$. The algorithm, Total_Transitivity(T), is outlined in Algorithm 1.
@@ 5.2 Proof of correctness
Similar lemma can be defined for modified-total transitivity.
@@
Let $v$ be a vertex of $T$, and $mttr(v,T)=t$. Then there exists a modified-total transitive partition of $T$, say $\{V_{1},V_{2},\ldots,V_{i}\}$, such that $v\in V_{i}$, for all $1\leq i\leq t$.
@@
Since $mttr(v,T)=t$, there is a modified-transitive partition $\pi=\{U_{1},U_{2},\ldots,U_{t}\}$ of $T$ such that $v\in U_{t}$. Similarly, we can show that there exists a modified-total transitive partition of $T$, say $\{V_{1},V_{2},\ldots,V_{i}\}$, such that $v\in V_{i}$, for all $1\leq i\leq t$. ∎
In the next theorem, we find the total transitive number or modified-total transitive number of a non-leaf and non-support vertex using the total transitive number or modified-total transitive number of its children.
@@
Let $v_{1},v_{2},\ldots,v_{k}$ be the children of $x$ in a rooted tree $T^{x}$, and for each $1\leq i\leq k$, $l_{i}$ be either $ttr^{r}(v_{i},T^{x})$ or $mttr^{r}(v_{i},T^{x})$. If $l_{s}$ is both $ttr^{r}(v_{s},T^{x})$ and $mttr^{r}(v_{s},T^{x})$ for some $s$, in that case we consider $l_{s}$ as $ttr^{r}(v_{s},T^{x})$ only. Furthermore, let $l_{j}^{\prime}=l_{j}$, if either $l_{j}=ttr^{r}(v_{j},T^{x})$ otherwise $l_{j}^{\prime}=l_{j}-1$ and $l_{1}^{\prime}\leq l_{2}^{\prime}\leq\ldots\leq l_{k}^{\prime}$. Assume $z$ be the largest integer such that there exists a subsequence of $\{l_{i}^{\prime}:1\leq i\leq k\}$, say, $(l_{i_{1}}^{\prime}\leq l_{i_{2}}^{\prime}\leq\ldots\leq l_{i_{z}}^{\prime})$ such that $l_{i_{p}}^{\prime}\geq p$, for all $1\leq p\leq z$. Also, assume $q\in\{i_{z}+1,i_{z}+2,\ldots,k\}$ such that $l_{q}$ is a $mttr^{r}(v_{q},T^{x})$ and $l_{q}=l_{q}^{\prime}+1\geq z+1$. Then, either total transitive number or total modified-transitive number of $x$ in the underlying tree $T$ is as follows.
@@
    If there exists $q\in\{i_{z}+1,i_{z}+2,\ldots,k\}$ such, that $l_{q}$ is a $mttr^{r}(v_{q},T^{x})$ and $l_{q}=l_{q}^{\prime}+1\geq z+1$. Then we have total transitive number of $x$ in the underlying tree, $T$, and it will be $z+1$, that is $ttr(x,T^{x})=1+z$.
@@
    If no such $q\in\{i_{z}+1,i_{z}+2,\ldots,k\}$ exits, such that $l_{q}$ is a $mttr^{r}(v_{q},T^{x})$ and $l_{q}=l_{q}^{\prime}+1\geq z+1$. Then we have modified-total transitive number of $x$ in the underlying tree, $T$, and it will be $z+1$, that is $mttr(x,T^{x})=1+z$.
@@
(a) First, consider the case when there exists $q\in\{i_{z}+1,i_{z}+2,\ldots,k\}$ such that $l_{q}$ is a $mttr^{r}(v_{q},T^{x})$ and $l_{q}=l_{q}^{\prime}+1\geq z+1$. Now, for each $j$, $1\leq j\leq z$, let us consider the subtrees $T_{v_{i_{j}}}^{x}$. It is also given that $ttr^{r}(v_{i_{j}},T^{x})=l_{i_{j}}^{\prime}$, for $j\in\{1,2,\ldots,z\}$. For all $1\leq p\leq z$, since $l_{i_{p}}^{\prime}\geq p$, by Lemma 25, we know that there exists a total transitive partition $\pi^{p}=\{V_{1}^{p},V_{2}^{p},\ldots,V_{p}^{p}\}$ of $T_{v_{i_{p}}}^{x}$, such that $v_{i_{p}}\in V_{p}^{p}$. Since $q$ exits and $mttr^{r}(v_{q},T^{x})=l_{q}=l_{q}^{\prime}+1\geq z+1$, by Lemma 26, we know that there exists a modified-total transitive partition $\pi^{q}=\{V_{1}^{q},V_{2}^{q},\ldots,V_{z+1}^{q}\}$ of $T_{v_{q}}^{x}$, such that $v_{q}\in V_{z+1}^{q}$. Let us consider a vertex partition $\pi=\{V_{1},V_{2},\ldots,V_{z},V_{z+1}\}$ of $T^{x}$ as follows: $V_{i}=(\displaystyle{\bigcup_{j=i}^{z}V_{i}^{j}})\cup V_{i}^{q}$, for $2\leq i\leq z$, $V_{z+1}=\{v_{q},x\}$ and every other vertices of $T^{x}$ are put in $V_{1}$. Since $x$ is not a leaf vertex or a support vertex, $\pi$ is a total transitive partition of $T^{x}$. Therefore, by the definition of total transitive number $ttr(x,T^{x})\geq 1+z$.
Next, we show that $ttr(x,T^{x})$ cannot be more than $1+z$. If possible, let $trr(x,T^{x})\geq 2+z$. Then, by Lemma 25, we have that there exists a transitive partition $\pi=\{V_{1},V_{2},\ldots,V_{2+z}\}$ such that $x\in V_{2+z}$. This implies that for each $1\leq i\leq 1+z$, $V_{i}$ contains a neighbour of $x$, say, $v_{i}$ such that the rooted total transitive number of $v_{i}$ is greater or equal to $i$, that is, $l_{i}=ttr^{r}(v_{i},T^{x})\geq i$. The set $\{l_{i}|1\leq i\leq 1+z\}$ forms a desired subsequence of $\{l_{i}:1\leq i\leq k\}$, contradicting the maximality of $z$. Hence, $ttr(x,T)=1+z$.
@@
(b) Let us assume no such $q\in\{i_{z}+1,i_{z}+2,\ldots,k\}$ exits, such that $l_{q}$ is a $mttr^{r}(v_{q},T^{x})$ and $l_{q}=l_{q}^{\prime}+1\geq z+1$. Now, for each $1\leq j\leq z$, let us consider the subtrees $T_{v_{i_{j}}}^{x}$. It is also given that $ttr^{r}(v_{i_{j}},T^{x})=l_{i_{j}}^{\prime}$, for $j\in\{1,2,\ldots,z\}$. For all $1\leq p\leq z$, since $l_{i_{p}}^{\prime}\geq p$, by Lemma 25, we know that there exists a total transitive partition $\pi^{p}=\{V_{1}^{p},V_{2}^{p},\ldots,V_{p}^{p}\}$ of $T_{v_{i_{p}}}^{x}$, such that $v_{i_{p}}\in V_{p}^{p}$. Let us consider the partition of $\pi=\{V_{1},V_{2},\ldots,V_{z},V_{z+1}\}$ of $T^{x}$ as follows: $V_{i}=\displaystyle{\bigcup_{j=i}^{z}V_{i}^{j}}$, for $2\leq i\leq z$, $V_{z+1}=\{x\}$ and every other vertices of $T^{x}$ are put in $V_{1}$. Since $x$ is not a leaf vertex or a support vertex, $\pi$ is a modified-total transitive partition of $T^{x}$. Therefore, $mttr(x,T)\geq 1+z$. Similarly, as above, we can show that $mttr(x,T)=1+z$. ∎
@@
Let $x$ be neither a leaf vertex nor a support vertex of a rooted tree $T^{c}$. Also, assume $v_{1},v_{2},\ldots,v_{k}$ are the children of $x$ in a rooted tree $T^{x}$, and for each $1\leq i\leq k$, $l_{i}$ be either $ttr^{r}(v_{i},T^{x})$ or $mttr^{r}(v_{i},T^{x})$. If $l_{s}$ is both $ttr^{r}(v_{s},T^{x})$ and $mttr^{r}(v_{s},T^{x})$ for some $s$, in that case we consider $l_{s}$ as $ttr^{r}(v_{s},T^{x})$ only. Furthermore, let $l_{j}^{\prime}=l_{j}$, if $l_{j}=ttr^{r}(v_{j},T^{x})$, otherwise, $l_{j}^{\prime}=l_{j}-1$ and $l_{1}^{\prime}\leq l_{2}^{\prime}\leq\ldots\leq l_{k}^{\prime}$. Now from Lemmas 25, 26 and Theorem 27, we have the following algorithm.
@@
Output: Either $ttr(x,T^{x})$ or $mttr(x,T^{x})$.
@@
2:Find $q\in\{i_{z}+1,i_{z}+2,\ldots,k\}$ such that $l_{q}$ is a $mttr^{r}(v_{q},T^{x})$ and $l_{q}=l_{q}^{\prime}+1\geq z+1$.

### v2

@@ 4 NP-completeness
5 Bipartite chain graphs
In this section, we design a linear-time algorithm for finding total transitivity of a given bipartite chain graph $G$. It is known that transitivity in a bipartite chain graph $G$ is $t+1$ if and only if $G$ contains either $K_{t,t}$ or $K_{t,t}\setminus\{e\}$ as an induced subgraph for the maximum value of such $t$ [18]. To design the algorithm for total transitivity, we first prove that total transitivity of a bipartite chain graph $G$ can be either $Tr(G)-2$ or $Tr(G)-1$, where $Tr(G)$ is transitivity of $G$.
Lemma 24.
Let $G=(X\cup Y,E)$ be a bipartite chain graph, and $\sigma_{X}=(x_{1},x_{2},\ldots,x_{n_{1}})$ and $\sigma_{Y}=(y_{1},y_{2},\ldots,y_{n_{2}})$ be the chain ordering of $G$. Also, assume that $Tr(G)=t+1$. Then the total transitivity of $G$, $Tr(G)-2\leq Tr_{t}(G)\leq Tr(G)-1$.
Proof.
Since $G$ is a bipartite chain graph and $Tr(G)=t+1$, $G$ contains either $K_{t,t}$ or $K_{t,t}\setminus\{e\}$ as an induced subgraph for the maximum value of such $t$ [18]. Moreover, from [18], we know that either $X_{t}\cup Y_{t}=V(K_{t,t})$ or $V(K_{t,t}\setminus\{e\})$ where $X_{t}=\{x_{1},x_{2},\ldots,x_{t}\}$ and $Y_{t}=\{y_{1},y_{2},\ldots,y_{t}\}$ and $\sigma_{X}=(x_{1},x_{2},\ldots,x_{m})$ and $\sigma_{Y}=(y_{1},y_{2},\ldots,y_{n})$ are the chain ordering of $G$. To prove that $Tr_{t}(G)\geq Tr(G)-2$, let us consider the following vertex partition $\pi=\{V_{1},V_{2},\ldots,V_{t},V_{t-1}\}$ as follows: $\{x_{1},y_{1}\}\subseteq V_{1}$, $V_{j}=\{x_{j},y_{j}\}$ for all $12\leq j\leq t-$ 1 and all the other vertices of $G$ in $V_{1}$ (See Figure 6).
Figure 6: Left represents total transitive partition of $G$ of size at least $Tr(G)-2$ when $G$ contains $K_{t,t}$, and right represents total transitive partition of $G$ of size at least $Tr(G)-2$ when $G$ contains $K_{t,t}\setminus\{e\}$
Clearly, in $\pi$, $V_{i}$ totally dominates $V_{j}$ for all $1\leq i\leq j\leq k$. Thus, $\pi$ is a total transitive partition of $G$ with size $t-1$, that is, $Tr(G)-2$. Therefore, $Tr_{t}(G)\geq Tr(G)-2$. To prove that $Tr_{t}(G)\leq Tr(G)-1$ for a bipartite chain graph, let us assume that $Tr_{t}(G)\geq Tr(G)$. In this case, by the definition of total transitivity, we have $Tr_{t}(G)=Tr(G)=t+1$. Let $\pi=\{V_{1},V_{2},\ldots,V_{t+1}\}$ be a total transitive partition of $G$ of size $t+1$. Since $|X_{t}|=t$, there exists at least one set in $\pi$, say $V_{r}$, such that $V_{r}$ does not contain any vertex from $X_{t}$. Since $\pi$ is a total transitive partition of $G$, each $V_{i}$ contains at least one vertex from $X$ and at least one vertex from $Y$.
Assume $x_{p}\in V_{r}$ for some $t+1\leq p\leq n_{1}$. Since $V_{r}$ totally dominates $V_{r}$, there exists $y_{\alpha}\in Y_{t}$ such that $y_{\alpha}\in V_{r}$. Again, $|Y_{t}|=t$ and $y_{\alpha}\in Y_{t}\cap V_{r}$, implying that there exists $V_{s}\in\pi$ with $s\neq t$, such that $V_{s}$ does not contain any vertex from $Y_{t}$. Similarly, as $V_{s}$ totally dominates $V_{s}$, there exists $x_{\beta}\in X_{t}$ such that $x_{\beta}\in V_{s}$. Therefore, we have $\{x_{\beta},y_{q}\}\subseteq V_{s}$ and $\{x_{p},y_{\alpha}\}\subseteq V_{r}$. Assume the case when $s<r$. Since $\pi$ is a total transitive partition, to totally dominate $x_{p}$, there must be a vertex from $Y_{t}$ in $V_{s}$, which contradicts the fact that $V_{s}$ does not contain a vertex from $Y_{t}$. Similarly, we can show a contradiction when $s>r$. Therefore, our assumption that $Tr_{t}(G)\geq Tr(G)$ is false. Hence, for a bipartite chain graph $G$, we have $Tr(G)-2\leq Tr_{t}(G)\leq Tr(G)-1$. ∎
Next, we characterize the bipartite chain graphs with total transitivity equal to $Tr(G)-1$.
Lemma 25.
Let $G=(X\cup Y,E)$ be a bipartite chain graph, and $\sigma_{X}=(x_{1},x_{2},\ldots,x_{n_{1}})$ and $\sigma_{Y}=(y_{1},y_{2},\ldots,y_{n_{2}})$ be the chain ordering of $G$. Also, assume that $Tr(G)=t+1$. Then $Tr_{t}(G)=Tr(G)-1$ if and only if $G$ contains $K_{t,t}$ as an induced subgraph for the maximum integer $t$.
Proof.
First, assume that $G$ contains $K_{t,t}$ as an induced subgraph for the maximum integer $t$. Now consider a vertex partition of $G$, say $\pi=\{V_{1},V_{2},\ldots,V_{t}\}$, such that $\{x_{1},y_{1}\}\subseteq V_{1}$, $V_{j}=\{x_{j},y_{j}\}$ for all $2\leq j\leq t$, and all the other vertices of $G$ in $V_{1}$. Since $G[X_{t}\cup Y_{t}]=K_{t,t}$, $V_{i}$ totally dominates $V_{j}$ for all $1\leq i\leq j\leq k$. Hence, $\pi$ is a total transitive partition of $G$. Now, by Lemma 24, we have $Tr_{t}(G)=Tr(G)-1$.
Conversely, let $Tr_{t}(G)=Tr(G)-1$. As $Tr(G)=t+1$, from the results in [18], we know that $G$ contains either $K_{t,t}$ or $K_{t,t}\setminus\{e\}$ as an induced subgraph for the maximum value of such $t$. If $G$ contains $K_{t,t}$ as an induced subgraph for the maximum value of such $t$, we are done. Assume the other part that is $G$ contains $K_{t,t}\setminus\{e\}$ as an induced subgraph for the maximum value of such $t$. We show that $Tr_{t}(G)=Tr(G)-2$, which contradicts our assumption. To show that $Tr_{t}(G)=Tr(G)-2$ when $G$ contains $K_{t,t}\setminus\{e\}$ as an induced subgraph for the maximum value of such $t$, we have the following claim.
Claim 26.
Let $G$ be a bipartite graph such that $G$ contains $K_{t,t}\setminus\{e\}$ as an induced subgraph for the maximum value of such $t$. Then $Tr_{t}(G)=Tr(G)-2$
Proof.
We show that $Tr_{t}(G)$ cannot be more than $Tr(G)-2$ for a graph $G$ when $G$ contains $K_{t,t}\setminus\{e\}$ as an induced subgraph for the maximum value of such $t$. If possible, assume $Tr_{t}(G)\geq Tr(G)-1$. From Lemma 24, we have $Tr_{t}(G)\leq Tr(G)-1$. So, in this case, $Tr_{t}(G)=Tr(G)-1=t$. Let $\pi=\{V_{1},V_{2},\ldots,V_{t}\}$ be a total transitive partition of $G$. Also, assume that $x_{t}\in V_{r}$ for some $r\in\{1,2,\ldots,t\}$. Since $\pi$ is a total transitive partition, for $x_{t}\in V_{r}$, there exists $y_{\alpha}\in\{y_{1},y_{2},\ldots,y_{t-1}\}$ such that $y_{\alpha}\in V_{r}$. Thus, there exists $V_{s}$, where $s\neq t$, such that $V_{s}$ does not contain any vertex from $\{y_{1},y_{2},\ldots,y_{t-1}\}$. Let $y_{\alpha^{\prime}}\in V_{s}$ for $\alpha^{\prime}\in\{t,t+1,\ldots,n_{2}\}$.
Now, if $r<s$, then to totally dominate $y_{\alpha^{\prime}}$, $V_{r}$ must contain a vertex from $\{x_{1},x_{2},\ldots,x_{t-1}\}$. Also, from the fact that $\pi$ is a total transitive partition, we have a vertex $x_{\beta}\in\{x_{1},x_{2},\ldots,x_{t-1}\}\cap V_{r}$. Therefore, in $\pi$, there exists $V_{p}$ for $p\neq s,r$ such that $V_{p}$ does not contain any vertex from $X_{t}$. Let $x_{\beta^{\prime}}\in V_{p}$ with $t+1\leq\alpha\leq n_{1}$. Clearly, $p>s$, otherwise, to totally dominate $y_{\alpha^{\prime}}$, there must be a vertex from $X_{t}$ in $V_{p}$. For this $x_{\beta^{\prime}}\in V_{p}$, there must be a vertex from $\{y_{1},y_{2},\ldots,y_{t-1}\}$ in $V_{p}$. So, there exists a set $V_{q}$ other than $V_{p}$, $V_{r}$, and $V_{s}$ in $\pi$, such that it does not contain any vertices from $\{y_{1},y_{2},\ldots,y_{t-1}\}$. Let us assume $y_{\alpha^{\prime\prime}}\in V_{q}$ and $\alpha^{\prime\prime}\in\{t,t+1,\ldots,n_{2}\}$. If $p>q$, to totally dominate $x_{\beta}$, there must be a vertex from $Y_{t}\setminus\{y_{t}\}$ in $V_{q}$. On the other hand, if $q>p$, then to totally dominate $y_{\alpha^{\prime\prime}}$, there must be a vertex from $X_{t}$ in $V_{p}$. Both cases are not possible. Similarly, we can show contradictions when $s<r$. Therefore, our assumption is false. Hence, $Tr_{t}(G)=Tr(G)-2$. ∎
As we assume $Tr_{t}(G)=Tr(G)-1$, from the above claim we can say that $G$ contains $K_{t,t}$ as an induced subgraph for the maximum value of such $t$. Hence we have the lemma. ∎
Based on Lemmas 24 and 25, we can design an algorithm for finding total transitivity of a given bipartite chain graph. In [18], authors showed that transitivity can be solved in linear time for a bipartite chain graph. In their algorithm they showed whether $G$ contains an induced $K_{t,t}$ or an induced $K_{t,t}\setminus\{e\}$. Hence, we have the following theorem.
Theorem 27.
The Maximum Total Transitivity Problem can be solved in linear time for bipartite chain graphs.
@@ 6 Polynomial-time algorithm for trees
In the Section 4, we have seen that Maximum Total Transitivity Decision Problem is NP-complete for bipartite graphs. In this section, we design a polynomial-time algorithm for finding the total transitivity of a given tree $T=(V,E)$. First, we give a comprehensive description of our proposed algorithm.
@@ 6.1 Description of the algorithm
To this end, we define another two parameters, namely the weak total transitive number and rooted weak total transitive number. The weak total transitive number of a vertex $v$ in $T$ is the maximum integer $p$ such that $v\in V_{p}$ in a weak total transitive partition $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$. Note that the maximum is taken over all weak total transitive partitions of $T$. We denote the weak total transitive number of a vertex $v$ in $T$ by $wttr(v,T)$. Finally, we define another parameter, namely the rooted weak total transitive number. The rooted weak total transitive number of $v$ in $T^{c}$ is the weak total transitive number of $v$ in the tree $T_{v}^{c}$, and it is denoted by $wttr^{r}(v,T^{c})$. Therefore, $wttr^{r}(v,T^{c})=wttr(v,T_{v}^{c})$.
Note that the value of the rooted total transitive number of a vertex is dependent on the rooted tree, whereas the total transitive number is independent of the rooted tree. Also, for the root vertex $c$, $ttr^{r}(c,T^{c})=ttr(c,T)$. We recursively compute the rooted total transitive number or rooted weak total transitive number of the vertices of $T^{c}$ in a bottom-up approach. First, we consider a vertex ordering $\sigma$, the reverse of the BFS ordering of $T^{c}$. For a leaf vertex $c_{i}$, we set $wttr^{r}(c_{i},T^{c})=1$ and for a support vertex $c_{j}$, $ttr^{r}(c_{j},T^{c})=wttr^{r}(c_{j},T^{c})=1$. For other non-leaf vertex $c_{i}$, which is not a support vertex, we call the function Total_Transitive_Number()()( ), which takes the rooted total transitive number or rooted weak total transitive number of children of $c_{i}$ in $T^{c}$ as input and returns the rooted total transitive number of $c_{i}$ in $T^{c}$ or rooted weak total transitive number $c_{i}$ in $T^{c}$ according to Theorem 30. At the end of the bottom-up approach, we have the rooted total transitive number of $c$ in $T^{c}$ or rooted weak total transitive number of $c$ in $T^{c}$, that is, $ttr^{r}(c,T^{c})$ or $wttr^{r}(c,T^{c})$, which is the same as the total transitive number of $c$ in $T$, that is, $ttr(c,T)$ or $ttr(c,T)+1$. After the bottom-up approach, we have the total transitive number of the root vertex $c$ and the rooted total transitive number or rooted weak total transitive number of every other vertex in $T^{c}$. Similarly, we compute the total transitive number of all vertices of $T$ by considering a rooted tree rooted at vertex $x$. Finally, we compute the maximum overall total transitive number of vertices in $T$ and return that value as the total transitivity of $T$. The algorithm, Total_Transitivity(T), is outlined in Algorithm 1.
@@ 6.2 Proof of correctness
Similar lemma can be defined for weak total transitivity.
@@
Let $v$ be a vertex of $T$, and $wttr(v,T)=t$. Then there exists a weak total transitive partition of $T$, say $\{V_{1},V_{2},\ldots,V_{i}\}$, such that $v\in V_{i}$, for all $1\leq i\leq t$.
@@
Since $wttr(v,T)=t$, there is a weak total transitive partition $\pi=\{U_{1},U_{2},\ldots,U_{t}\}$ of $T$ such that $v\in U_{t}$. Similarly, we can show that there exists a weak total transitive partition of $T$, say $\{V_{1},V_{2},\ldots,V_{i}\}$, such that $v\in V_{i}$, for all $1\leq i\leq t$. ∎
In the next theorem, we find the total transitive number or weak total transitive number of a non-leaf and non-support vertex using the total transitive number or weak total transitive number of its children.
@@
Let $v_{1},v_{2},\ldots,v_{k}$ denote the children of $x$ in a rooted tree $T^{x}$, and for each $1\leq i\leq k$, let $l_{i}$ be either $ttr^{r}(v_{i},T^{x})$ or $wttr^{r}(v_{i},T^{x})$. If $l_{s}$ is both $ttr^{r}(v_{s},T^{x})$ and $wttr^{r}(v_{s},T^{x})$ for some $s$, we will regard $l_{s}$ solely as $ttr^{r}(v_{s},T^{x})$. Additionally, let $l_{j}^{\prime}=l_{j}$ if $l_{j}=ttr^{r}(v_{j},T^{x})$; otherwise, let $l_{j}^{\prime}=l_{j}-1$, ensuring that $l_{1}^{\prime}\leq l_{2}^{\prime}\leq\ldots\leq l_{k}^{\prime}$. Let $z$ denote the largest integer for which there exists a subsequence of $\{l_{i}^{\prime}:1\leq i\leq k\}$, denoted as $(l_{i_{1}}^{\prime}\leq l_{i_{2}}^{\prime}\leq\ldots\leq l_{i_{z}}^{\prime})$, satisfying the condition that $l_{i_{p}}^{\prime}\geq p$ for all $1\leq p\leq z$. Moreover, let $q\in\{i_{z}+1,i_{z}+2,\ldots,k\}$ such that $l_{q}$ is a $wttr^{r}(v_{q},T^{x})$ and $l_{q}=l_{q}^{\prime}+1\geq z+1$. The total transitive number or total weak transitive number of $x$ is defined in the underlying tree $T$ as follows.
@@
    If there exists $q\in\{i_{z}+1,i_{z}+2,\ldots,k\}$ such that $l_{q}$ is a $wttr^{r}(v_{q},T^{x})$ and $l_{q}=l_{q}^{\prime}+1\geq z+1$. Then we have total transitive number of $x$ in the underlying tree, $T$, and it will be $z+1$, that is, $ttr(x,T^{x})=1+z$.
@@
    If no such $q\in\{i_{z}+1,i_{z}+2,\ldots,k\}$ exists, such that $l_{q}$ is a $wttr^{r}(v_{q},T^{x})$ and $l_{q}=l_{q}^{\prime}+1\geq z+1$. Then we have weak total transitive number of $x$ in the underlying tree, $T$, and it will be $z+1$, that is, $wttr(x,T^{x})=1+z$.
@@
(a)First, consider the case when there exists $q\in\{i_{z}+1,i_{z}+2,\ldots,k\}$ such that $l_{q}$ is a $wttr^{r}(v_{q},T^{x})$ and $l_{q}=l_{q}^{\prime}+1\geq z+1$. Now, for each $j$, $1\leq j\leq z$, let us consider the subtrees $T_{v_{i_{j}}}^{x}$. It is also given that $ttr^{r}(v_{i_{j}},T^{x})=l_{i_{j}}^{\prime}$, for $j\in\{1,2,\ldots,z\}$. For all $1\leq p\leq z$, since $l_{i_{p}}^{\prime}\geq p$, by Lemma 28, we know that there exists a total transitive partition $\pi^{p}=\{V_{1}^{p},V_{2}^{p},\ldots,V_{p}^{p}\}$ of $T_{v_{i_{p}}}^{x}$, such that $v_{i_{p}}\in V_{p}^{p}$.
@@
Since $q$ exists and $wttr^{r}(v_{q},T^{x})=l_{q}=l_{q}^{\prime}+1\geq z+1$, by Lemma 29, we know that there exists a weak total transitive partition $\pi^{q}=\{V_{1}^{q},V_{2}^{q},\ldots,V_{z+1}^{q}\}$ of $T_{v_{q}}^{x}$, such that $v_{q}\in V_{z+1}^{q}$. Let us consider a vertex partition $\pi=\{V_{1},V_{2},\ldots,V_{z},V_{z+1}\}$ of $T^{x}$ as follows: $V_{i}=\left(\bigcup_{j=i}^{z}V_{i}^{j}\right)\cup V_{i}^{q}$, for $2\leq i\leq z$, $V_{z+1}=\{v_{q},x\}$, and all other vertices of $T^{x}$ are placed in $V_{1}$. Since $x$ is neither a leaf vertex nor a support vertex, $\pi$ is a total transitive partition of $T^{x}$. Therefore, by the definition of the total transitive number, $ttr(x,T^{x})\geq 1+z$.
Next, we show that $ttr(x,T^{x})$ cannot be more than $1+z$. If possible, suppose $ttr(x,T^{x})\geq 2+z$. Then, by Lemma 28, there exists a total transitive partition $\pi=\{V_{1},V_{2},\ldots,V_{2+z}\}$ such that $x\in V_{2+z}$. This implies that for each $1\leq i\leq 1+z$, $V_{i}$ contains a neighbour of $x$, say $v_{i}$, such that the rooted total transitive number of $v_{i}$ is greater than or equal to $i$, i.e., $l_{i}=ttr^{r}(v_{i},T^{x})\geq i$. The set $\{l_{i}\mid 1\leq i\leq 1+z\}$ forms the desired subsequence of $\{l_{i}\mid 1\leq i\leq k\}$, contradicting the maximality of $z$. Hence, $ttr(x,T^{x})=1+z$.
(b) Let us assume no such $q\in\{i_{z}+1,i_{z}+2,\ldots,k\}$ exits, such that $l_{q}$ is a $wttr^{r}(v_{q},T^{x})$ and $l_{q}=l_{q}^{\prime}+1\geq z+1$. Now, for each $1\leq j\leq z$, let us consider the subtrees $T_{v_{i_{j}}}^{x}$. It is also given that $ttr^{r}(v_{i_{j}},T^{x})=l_{i_{j}}^{\prime}$, for $j\in\{1,2,\ldots,z\}$. For all $1\leq p\leq z$, since $l_{i_{p}}^{\prime}\geq p$, by Lemma 28, we know that there exists a total transitive partition $\pi^{p}=\{V_{1}^{p},V_{2}^{p},\ldots,V_{p}^{p}\}$ of $T_{v_{i_{p}}}^{x}$, such that $v_{i_{p}}\in V_{p}^{p}$. Let us consider the partition of $\pi=\{V_{1},V_{2},\ldots,V_{z},V_{z+1}\}$ of $T^{x}$ as follows: $V_{i}=\displaystyle{\bigcup_{j=i}^{z}V_{i}^{j}}$, for $2\leq i\leq z$, $V_{z+1}=\{x\}$ and every other vertices of $T^{x}$ are put in $V_{1}$. Since $x$ is not a leaf vertex or a support vertex, $\pi$ is a weak total transitive partition of $T^{x}$. Therefore, $wttr(x,T)\geq 1+z$. Similarly, as above, we can show that $wttr(x,T)=1+z$. ∎
@@
Let $x$ be neither a leaf vertex nor a support vertex of a rooted tree $T^{c}$. Also, assume $v_{1},v_{2},\ldots,v_{k}$ are the children of $x$ in a rooted tree $T^{x}$, and for each $1\leq i\leq k$, $l_{i}$ be either $ttr^{r}(v_{i},T^{x})$ or $wttr^{r}(v_{i},T^{x})$. If $l_{s}$ is both $ttr^{r}(v_{s},T^{x})$ and $wttr^{r}(v_{s},T^{x})$ for some $s$, in that case we consider $l_{s}$ as $ttr^{r}(v_{s},T^{x})$ only. Furthermore, let $l_{j}^{\prime}=l_{j}$, if $l_{j}=ttr^{r}(v_{j},T^{x})$, otherwise, $l_{j}^{\prime}=l_{j}-1$ and $l_{1}^{\prime}\leq l_{2}^{\prime}\leq\ldots\leq l_{k}^{\prime}$. Now from Lemmas 28, 29 and Theorem 30, we have the following algorithm.
@@
Output: Either $ttr(x,T^{x})$ or $wttr(x,T^{x})$.
@@
2:Find $q\in\{i_{z}+1,i_{z}+2,\ldots,k\}$ such that $l_{q}$ is a $wttr^{r}(v_{q},T^{x})$ and $l_{q}=l_{q}^{\prime}+1\geq z+1$.